### v1

# “Talking Past Each Other”: Issue Ownership and Microtargeting in Swiss Online Political Ads

Abstract
Switzerland’s unique system of direct democracy, characterized by frequent popular referenda, provides a critical context for studying the impact of online political advertising beyond standard electoral cycles. This paper presents a large-scale, data-driven analysis of 40k political ads published on Facebook and Instagram in Switzerland between 2021 and 2025. Despite a voting population of only 5.6 million, the ad campaigns were significant in scale, costing CHF 4.5 million and achieving 560 million impressions. This study shows that political ads are used not only for federal elections, but also to influence referenda, where greater exposure to “pro-Yes” advertising correlates significantly with approval outcomes. The analysis of microtargeting reveals distinct partisan strategies: centrist and right-wing parties predominantly target older men, whereas left-wing parties focus on young women. Furthermore, significant region-specific demographic variations are observed even within the same party, reflecting Switzerland’s strong territorial divisions. Regarding content, a clear pattern of “talking past each other” is identified: in line with issue ownership theory, parties avoid direct debate on shared issues, preferring to promote exclusively owned topics. Finally, it is demonstrated that these strategies are so distinct that an ad’s author can be predicted using a machine learning model trained exclusively on its audience and topic features. This study sheds light on how microtargeting and issue divergence on social platforms may fragment the public sphere and bypass traditional democratic deliberation.
1 INTRODUCTION
The efficacy of direct democracy is often questioned in academic literature, with some analyses finding it to be ineffective or even counterproductive in practice. Critics often highlight the potential for a “tyranny of the majority”, the influence of wealthy special interest groups, and the risk posed by citizen incompetence regarding complex policy issues [warner1995direct, gerber2011populist]. However, Switzerland is often cited as an exception to this criticism, with its political system considered a model of success [BrunoFrey]. This Swiss model is not a simple and universally applicable template; rather, it has developed within a specific and unique historical and institutional setting. It evolved "bottom-up" from cantonal traditions, not being imposed by a central authority [Linder2010].
Swiss citizens vote on national issues approximately four times per year, participating in far more referendums and initiatives than citizens of any other country [Serdült2014]. This high degree of participation is argued to have numerous positive social and economic effects. These include empowering citizens, making procedures more appealing and imposing fiscal discipline on public spending [BrunoFrey, STUTZERALOISBRUNO]. Further research connects direct democratic rights directly to improved economic performance [FeldLarsSaviozMarcel] and higher levels of subjective well-being, or happiness, among citizens [FreyBrunoStutzerAlois2000].
Even within the Swiss context, the positive portrayal of direct democracy has its detractors. Some scholars have highlighted problematic aspects, including “voting fatigue” caused by the large number of decisions, the complexity of certain ballot measures, and the system’s ability to impede social and political reforms [BLONDELJEAN]. Furthermore, the “double majority” rule, which requires a majority of both citizens and cantons, has been criticized for giving disproportionate power to small, rural cantons. This rule enables controversial initiatives to pass despite lacking broad popular support in major population centers [Papadopoulos2001].
Beyond these long-standing debates, the proliferation of digital media introduces a new and critical challenge to direct democracy. In fact, online political advertising has long been considered controversial and potentially capable of undermining democratic processes. This idea is based on studies highlighting its effectiveness in manipulating voters, fueling cynicism, and polarizing the electorate [JamiesonKathleenHall]. The advent of microtargeting, which uses vast datasets to tailor persuasive messages to narrow, psychologically profiled segments of the electorate, significantly amplifies these threats and introduces new challenges [ZuiderveenBorgesiusFrederik]. In fact, personalized political ads are found to be more effective at persuasion than generic, non-personalized messages [SimchonEdwards, TappinWittenberg].
Active democracy is founded on the idea that citizens are well-informed and capable of debating and deciding on complex policy issues as part of shared public deliberation [BLONDELJEAN]. If microtargeting can effectively sway support with personalized, often emotional, and non-public messages, it threatens to bypass this traditional public sphere. Could new forms of digital propaganda potentially hinder effective democratic participation? What happens when microtargeting and opaque online political ad campaigns are exploited in a country that prides itself on direct democracy as a source of national identity?
These questions require a data-driven approach to assess the normative impact and understand the scale and scope of this phenomenon. This paper aims to present a comprehensive analysis of political advertising on social media in Switzerland, providing an empirical foundation for a critical debate. The main research questions are the following:
  • RQ1 Microtargeting
    a: Do parties target specific users based on demographics?
    b: Are there regional-specific demographics?
  • RQ2 Referenda
    a: Are ads used intensively during popular referenda?
    b: To what extent can ads influence referendum outcomes?
  • RQ3 Federal elections
    a: Which issues do the parties promote the most, and how?
    b: Is the content of an ad a reliable predictor of its author?
These research questions are addressed by collecting and analyzing 40,000 political ads published on Facebook and Instagram from 2021 to 2025. Through an analysis of demographic microtargeting and textual content, this study highlights the differences and similarities among Switzerland’s major political parties and linguistic regions. This is the first research in the literature to employ statistical models and techniques to examine such a large volume of online political ads published in Switzerland.
2 BACKGROUND AND RELATED WORK
2.1 Swiss Context
In order to understand the deep cultural roots of direct democracy in Swiss institutions and society, it is necessary to start with the very foundations of the Swiss Confederation. The modern federal state was established in 1848 as a direct result of a brief civil war, known as the Sonderbund War. The new state faced the challenges of integrating minorities and managing the country’s deep “cross-cutting cleavages” [Linder2010]. Switzerland was and remains a mosaic:
Linguistic. Since unification, there have been four national languages: German, French, Italian, and Romansh. In 2023, 61.4% of the population is native German speaker; 22.6% are native French speakers; 7.7% are native Italian speakers; and 0.5% are native Romansh speakers [FSOLang].
Religious. According to the Swiss National Survey [FSOrelsurvey], 35.6% of the population was religiously unaffiliated in 2023. The remaining population was divided as follows: 30.7% Catholic, 19.5% Swiss Protestant, 5.8% other Christian denominations, and 6% Muslim.
Geographic. Although rural municipalities cover 57% of the country’s total area, in 2025 they are home to only 35% of the population. In contrast, urban areas account for just 17% of the total area but are home to 65% of the population and 76% of the workforce [FSOGeo].
Economic. Cantons with major urban centers, such as Zurich and Basel-Stadt, have an exceptionally high economic output compared to other cantons. For instance, the nominal GDP per capita in Basel-Stadt was about CHF 210,000 in 2022, compared to the Swiss national average of around CHF 90,000 and the canton of Uri’s average of CHF 58,000 [FSO_GDP].
The federal state’s guarantee of municipal and cantonal autonomy with regard to taxation, education, and infrastructure has played an important role in preserving cohesion [Sager_Zollinger_2011]. Alongside this federal structure, the system of direct democracy has been the most important political compromise made to maintain cohesion in diversity since the unification in 1848. For this reason, Swiss direct democracy has often been interpreted by scholars as a mechanism for minority protection and power-sharing, specifically designed to integrate the country’s diverse segments [Linder2010, Disset2010]. The legal instruments available to citizens and enshrined in the Swiss Federal Constitution include the compulsory referendum, non-compulsory (facultative) referendum, and the popular initiative [Kriesi2005, Papadopoulos2001].
Swiss direct democracy operates within a historically highly fragmented, multipolar party system that reflects the nation’s underlying divisions. In 2025, the Federal Assembly remains highly multipolar, with no single party holding a majority. Following the 2023 federal elections, the national-conservative SVP (Swiss People’s Party) affirmed its position as the largest party with 27.9% of the vote. The center-left SP (Social Democratic Party) followed at 18.3%. The political center is split between the liberal, center-right FDP (FDP.The Liberals) at 14.3% and the Christian democratic Die Mitte (The Center) at 14.1%. The environmentalist bloc, which saw historic gains in 2019, is divided between the left-leaning GRÜNE (Green Party) at 9.8% and the centrist GLP (Green Liberal Party) at 7.6% [FSO_Elections2023]. Finally, it should be mentioned that in Switzerland, in 2025, out of a population of approximately 9.1 million inhabitants, only about 5.6 million are eligible to vote.
2.2 Online political advertising
The influence of social media platforms on political opinion formation and voting behavior remains a topic of debate in the scientific literature. For instance, large-scale experiments have demonstrated that algorithmically delivered social messages can measurably increase voter turnout [bond201261]. It has been proven that curated search rankings can also potentially shift the preferences of undecided voters [Epstein]. Other studies have examined the impact of online communication on politically sensitive topics. For example, Papakyriakopoulos4 investigated the role of demagoguery in civic engagement on Reddit, while LeeYu-Hao examined participation in online activism, or “slacktivism”, in donation campaigns. Memes and other new forms of communication have also been studied as a mean of political propaganda on social media [WilliamsNagel, Mejova, Leon].
This online political context intersects with the underlying business model of social media platforms, which is predominantly based on advertising [ANDERSON201541]. The economic logic incentivizes the continuous algorithmic optimization of content delivery systems to maximize user engagement and, consequently, advertising revenue [O’Reilly_Strauss_Mazzucato_2024, RanDezhi]. The exploitation of these highly optimized, profit-driven systems for the dissemination of political messages is, on the one hand, a source of contemporary public debate and academic research [DommettKatharine, AliSapiezynski]. On the other hand, the effectiveness of recommendation systems has also made online ads much more effective [ChuVliegenthart] and, as a result, led to a significant increase in their use [Sosnovik]. In this context, Meta’s platforms (Facebook and Instagram) stand out as among the most effective in influencing public opinion [KarpfDavid]. madrigal2017facebook argues that although Facebook’s potential impact on elections had been evident for years, it reached its peak during the 2016 presidential election, with \$1.4 billion spent on online ads during the campaign. However, hundreds of millions of dollars of dark ads paid for by foreign governments to influence elections should be added to this number [madrigal2017facebook]. In this regard, it has been proven that Russia attempted to manipulate the 2016 U.S. presidential election by running political ads on Facebook [entous2017russian, EmilioFerrara]. MatiasNathan designed a software-supported approach for auditing, which uses coordinated volunteers to analyze political advertising policies enacted by Facebook and Google during the 2018 midterm U.S. election. Other studies on Facebook political ads include anti-vaccination ads during COVID-19 [MejovaKalimeri], populist advertising for the European Parliamentary election in 2019 [CapozziThin], Italian political propaganda on migrants issue [capozzi2021clandestino, 10.1007/978-3-030-60975-7_4], or far-right misinformation spread by the populist Spanish party VOX [CanoLorena]. Many studies examined the spread and impact of political ads in Switzerland, with the earliest focusing on traditional newspapers [GerberBühlmann, ZumofenGuillaume]. Other studies have focused on referendum campaigns shifting from newspapers to social media [FischerMichaela], or politicians’ use of social media [ReveilhacMorselli].
3 DATA
This study relies on data from two main sources: the Swiss Federal Statistical Office [^1] https://abstimmungen.admin.ch/en/overview and the Meta Ad Library API [^2] https://www.facebook.com/ads/library/api/. The Federal Statistical Office provides detailed statistics related to referenda, including results by canton, participation percentages, and full descriptions of referendum questions. All data collected and annotated [^3] Zenodo anonymized URL and the code written [^4] GitHub anonymized URL are publicly available.
3.1 Meta Ad Library
The Meta Ad Library API discloses advertisements about social issues, election, or politics that were delivered on Facebook or Instagram during the past seven years. Although the API allows for specifying search parameters, such as limiting results to a specific country, all queries must contain keywords or author names. The following information can be extracted for each advertisement: the publication date; the end time of the campaign; the textual content; and the number of impressions (provided as an interval). The term “impressions” refers to non-unique visualizations. It is also possible to access the campaign expenditure (again presented as an interval) and the audience’s demographic profile, specifically their gender (male, female, or unknown) and age (categorized into seven buckets). In line with existing literature [capozzi2021clandestino], the midpoint of the interval’s boundaries is computed for values presented as ranges (expenditure and impressions). For unbounded ranges, only the defined boundary is used.
The collection is initiated using a broad set of election-related keywords (see Appendix Table A1) in German, French, and Italian due to the API’s requirement for keyword- or author-based searches. This process yielded an initial corpus of 40,044 ads (22,494 from German keywords, 10,794 from French, and 6,756 from Italian) published between January 1, 2021, and September 30, 2025. The API’s country filter includes advertisements displayed in Switzerland, even if they are intended for a different audience. To create a Swiss-centric dataset, ads are further filtered to include only those for which more than 50% of impressions come from users based in Switzerland. This reduces the dataset to 34,559 ads from 2,246 unique pages, which received an estimated 560 million impressions and cost over CHF 4.5 million.
Although the API provides a page name (username) for each ad, this identifier is frequently ambiguous or does not directly correspond to a particular political party, even when the content of the page clearly promotes a party’s agenda. As shown in Figure 1, the 2023 federal elections represent the most significant activity peak in the dataset. For this reason, the 1,224 pages that published at least one ad in the 30 days preceding the federal election are manually labeled to establish whether they are related to a specific party. This process identified 943 pages that could be directly attributed to a political party, while the remaining 281 pages belong to entities such as media organizations (newspapers and daily newspapers), NGOs, neighborhood committees, and non-partisan referendum groups. Of the 943 political pages, 17 major parties and smaller entities (e.g., municipal parties and civic lists) are identified. However, only six parties — SVP, SP, FDP, Die Mitte, GRÜNE, and GLP — play a dominant role in the advertising landscape, accounting for 95% of all impressions. Collectively, these parties also secured 91.84% of the votes in the 2023 federal elections. Tables  A2,  A3, and  A4 provide detailed statistics on impressions, volume, and ad expenditure broken down by party and language.
3.2 Content Annotation
The filtered dataset of 34,559 ads is cleaned to remove entries with blank or duplicate textual content, resulting in 13,324 unique ads. The paper presents two analyses of the textual content: one focused on referenda, and the other on the 2023 federal elections.
3.2.1 Referenda
The first step is to identify which advertisements relate to one of the 42 referenda held between January 2021 and October 2025. The process begins by filtering the ads to include only those containing at least 10 words and published within 30 days of the referendum. This procedure results in 9,118 ads. The second step involves prompting GPT-4o with the text of the ad campaign to establish whether it is related to a referendum and, if so, what stance the campaign is taking. In addition, the model is provided with the official title and description of each referendum, as published by the Federal Statistical Office. The model checkpoint is “gpt-4o-2024-08-06”, and the API queries are executed on October 8, 2025. The task is structured as a few-shot query, and the output for each ad is a Boolean parameter indicating whether the ad is relevant to the referendum. If an ad is deemed relevant, the output includes also the referendum’s title and the ad’s stance (Yes, No, or Neutral). Manual label verification is performed on a random sample of 100 ads, and this verification revealed an accuracy rate of 100%.
3.2.2 Federal Election
The second analysis focuses on the 2023 federal elections. Unlike a referendum, which addresses one or a few political issues, federal elections involve political parties promoting broader agendas. Therefore, it is necessary to not only identify election-relevant ads but also determine the specific topic(s) they cover. Consistent with the previous task, only ads with at least 10 words published within 30 days before the federal election are selected. The model checkpoint and the API query date are the same of the previous task. The model’s output includes a Boolean parameter identifying whether the ad is relevant to the federal election, as well as up to three English keywords related to the ad’s topic. This initial extraction process yielded a large set of 739 unique keywords, which require consolidation into fewer categories. To this end, each keyword is encoded into a semantic vector using OpenAI’s state-of-the-art text-embedding-3-large model. These vectors are then clustered into 25 topics using the K-Means clustering algorithm [mcqueen1967some]. To ensure reliability, qualitative validation is performed on a random subset of 100 ads. In this review, human annotation agreed with model-generated keywords in 96% of the cases.
4 RESULTS
Figure 1: Seven-day rolling average of ad impressions. The inset provides a daily view, disaggregated by language, for the 30-day period preceding the federal election.
4.1 Microtargeting
Figure 2 compares each party’s audience demographics to the “party complementary” profile, which consists of all advertisements not associated with that party. Since RQ1.a does not concern regional differences, ads from all languages are combined for this analysis. Gender disparities are quantified using odds ratios (OR), which measure the increased likelihood that an ad impression will be viewed by a man. An OR of 1 represents perfect gender balance; an OR less than 1 indicates a female bias; and an OR greater than 1 indicates a male bias. For age, the divergence between distributions is measured using the Wasserstein distance (WD), which indicates the mean difference in years for a party versus the complementary. Overall, centrist and right-wing parties are more oriented towards a male audience, particularly the SVP ($36\%$). On the other hand, the GRÜNE and the SP are more female-oriented, with the GRÜNE at $27\%$ and the SP at $13\%$. In terms of age reach, the SVP and, to a lesser extent, Die Mitte tend to attract an older audience, whereas the GRÜNE and the GLP primarily attract people in the 25–34 age group. Furthermore, the GRÜNE is the only party that extensively reaches the 18–24 age group, yet it underrepresents the over-45 age group more than any other party.
Figure 2: Comparison of the demographic reach by age and gender of each party’s ads versus all other ads. Gender is measured using male-to-female odds ratios, and age using Wasserstein distances.
To answer RQ1.b and evaluate the cantonal differences, the demographic distributions can be stratified by language. Figure 4 shows the male-to-female odds ratio within each party’s audience by language. Each cell (party/language) of the heatmap is computed with bootstrapping. For each of the 100 iterations, the OR is computed on a sample of the party’s ads and the complement’s ads with replacement. The mean and 95% confidence interval are computed at the end of the process from the bootstrap distribution. While center- and right-leaning parties attract male-dominated audiences in all language regions, the SVP’s audience in Ticino, the only predominantly Italian-speaking canton, is the most male-oriented. Other significant regional differences exist. For instance, the GLP and Die Mitte are gender-balanced in the French-speaking cantons and slightly male-oriented in the German-speaking cantons ($15\%$ and $20\%$, respectively). However, they are much more male-oriented in Ticino ($40\%$ and $28\%$, respectively). Further to the left on the political spectrum, the situation changes dramatically. For the SP, Ticino is again the least gender-balanced canton, but this time it favors a more female audience ($32\%$). In German-speaking cantons, the SP reaches a more female audience ($13\%$), while in French-speaking cantons, the audience is gender balanced. On the other hand the GLP is more gender unbalanced in French-speaking cantons ($34\%$ women).
Figure 3: Language-specific age distributions (bootstrapped mean and SD) for each party’s ads compared to its complementary ad set. The Wasserstein Distance measures the divergence between the two.
Figure 3 shows the age distribution of the audience reached by the parties, compared to the party complementaries, broken down by language group. The shaded area represents the standard deviation of the audience reached for each age group. The distribution is build using a bootstrapping procedure similar to that described for Figure 4. Significant differences between parties within the same canton can be observed, as well as regional variations within parties. While the SVP focuses on an audience over 45, especially in Ticino, Die Mitte focuses on an older audience in French-speaking cantons only. The SP, in German-speaking cantons, has a WD of almost 0 (no differences with the complementary party’s audience), while in French-speaking cantons, it targets a much more over-45 audience. In Ticino, however, it does the opposite, underrepresenting the over-35 audience and focusing on the 18-24 age range. The GLP, on the other hand, focuses heavily on the 25-34 age range in French-speaking cantons, while in Ticino, the reach is broader (25-45). With the exception of the GRÜNE and the SVP, parties in the German-speaking cantons show fewer differences compared to their counterparts. In summary, parties differ in their demographic reach (RQ1.a), with evident internal variations emerging across different linguistic regions (RQ1.b).
Figure 4: Male-to-female odds ratios for each party, relative to its complement, by language. Values are bootstrapped means, and 1 means perfect gender balance. An asterisk indicates that the 95% confidence interval does not include 1.0.
4.2 Referenda
Figure 1 shows the seven-day moving average of total political ad views. The gray vertical lines represent referendums, and the inset shows the total number of views per language in the 30 days before the federal elections of October 22, 2023. Clearly, this event caused the most significant increase in views. At first glance, it is instead difficult to answer RQ2.a and determine whether advertising activity increases before the 14 referendum dates. As preliminary step, a Wilcoxon signed-rank test [Wilcoxon] is applied to compare ad views and spending in the 30 days before and after each referendum. For completeness, this analysis is conducted separately for German, French, and Italian ads. The results of all the tests (shown in Table  A5) confirm a significant increase in ad views and spending in the 30 days before a referendum ($p<0.001$ in all cases).
Although the Wilcoxon signed-rank test confirms a difference, it does not account for confounding factors, such as seasonality or other temporal patterns. The true impact of the referenda can be isolated by fitting a Generalized Additive Model (GAM) to predict the daily spending and impressions. This model includes smooth functions for days until the referendum, the day of the year, and the trend, as well as a factor for the day of the week. Figure 5 shows the partial effects of the proximity of the referendum on the log-transformed expenditure and impressions. The results are consistent across all language groups and suggest a significant increase in ad spending and impressions 30 days before the vote. The 95% confidence intervals remain entirely above the baseline during this pre-referendum period, and then collapse immediately after the vote at $t=0$. These results confirm that ads are used intensively during the referendum campaign period.
Figure 5: Partial effect of days to referendum (t=0) on log-transformed Expenditure (blue) and Impressions (orange), shown by language and aggregate. Shaded areas represent 95% confidence intervals.
These findings lead to the next research question (RQ2.b): Given the increase in ad activity before referenda, is it possible to explore and measure the influence of referendum ads on the outcome of the vote? Figure 6 shows the distribution of all ad impressions supporting or opposing the referenda in relation to their outcomes. This figure suggests that there is a relationship between the number of impressions per side and referendum outcomes. When a referendum is rejected, the number of views of ads opposing the referendum looks to be higher. In contrast, if the outcome of a referendum is positive, the number of views of ads supporting it tends to be higher. Correlation analysis shows a moderate positive association between the ratio of Yes to No ad impressions and referendum approval outcome (Spearman $\rho=0.39$, $p=0.043$). This indicates that referendums with relatively more Yes-side advertising tended to be approved more often. Then, a Mann-Whitney U non-parametric test is used to independently compare the “Yes” and “No” impression volumes. The results suggest that approved referenda have a significantly higher volume of “Yes” impressions compared to rejected referenda ($U=146.0$, $p=0.0135$). In contrast, the volume of “No” impressions is not a statistically significant predictor of a referendum rejection ($U=117.0$, $p=0.1906$). In other words, greater exposure to pro-Yes advertising is associated with a higher probability of referendum approval, whereas exposure to pro-No advertising does not show a systematic link to rejection outcomes. It is important to note that this is a correlation and does not prove causation. In addition, popularity can act as a confounding factor: popular referendums are more likely to attract higher ad spending for the “Yes” side and are also more likely to pass. This is because “Yes” campaigns inherently face an uphill battle against the well-documented “status quo bias”, in which uncertain voters tend to vote “No” to avoid uncertainty [Bowler1998]. A “Yes” campaign must first have strong latent public support to overcome this bias and succeed. Thus, ad volume may indicate a well-funded campaign with this support rather than being an independent driver of the outcome [Gerber1998].
Figure 6: Distribution of ad impressions for pro-Yes and pro-No campaigns, grouped by referendum outcome.
Figure 7: Share of impressions per topic by party. Only the top ten topics for each party are represented.
4.3 Political issues
The third research question focuses on the federal elections in October 2023 and the issues promoted in ad campaigns. As described in 3.2.2, each ad published in the month before the federal elections is labelled with a relevance flag and with up to 3 keywords related to the topics discussed in the ad. Thanks to this data-processing procedure, it is possible to investigate whether parties discuss different issues during federal elections or promote different views on the same issues (RQ3.a).
Figure 7 shows the share of impressions by party for the top ten topics by impressions. Most issues are promoted exclusively by one party or by a few parties that are close to each other on the political spectrum. For example, the issues of “Business & Regulation” and “Taxation & Public Finance” are almost only represented by the SVP, the FDP, and Die Mitte. Most of the impressions on the issue of “Civil Liberties & Rights” (which includes topics such as abortion, drugs liberalization, and freedom of speech) are shared by the SVP and the FDP; on the other hand, most of the impressions on the issue “Agriculture & Food Security” are shared by the SVP and Die Mitte. Instead, issues related to “Pensions & Retirement”, “Climate & Environment”, “Energy & Policy”, and “Foreign Relations” (which covers topics such as EU relations, NATO, or the war in Ukraine) are shared almost exclusively by the SP and the GRÜNE. They also focus heavily on “Family & Youth Policy”, “Gender & LGBTQ+ Rights”, “Social Justice & Equality”, and “Political Spectrum” (including topics such as progressive policies and liberalism). However, other issues transcend traditional political divides, receiving support from parties that would otherwise be considered ideologically distant. These include “Education System” (shared by the SVP, the FDP, and the SP), “Infrastructure & Mobility” (shared by the GRÜNE and the SVP), Housing & Rent” (shared by the SP, the GRÜNE and the FDP), “Healthcare System” (shared by the SP, the GRÜNE and Die Mitte), and “Economy & labour” (shared by all parties except the GLP and Die Mitte).
Finally, some issues are pushed by a single party. For instance, “National Security” and “Immigration & Asylum” are cornerstones of the SVP’s agenda but are not promoted by other parties. The issue of “Media & Information” (which includes topics such as cancel culture, disinformation, and woke culture) is almost exclusively promoted by the center-right FDP party. The issue of “Urban & Regional Development” is addressed almost exclusively by the SP. It is also worth mentioning that the more centrist parties, GLP and Die Mitte, have the lowest share values. It is reasonable to hypothesize that this is because they address a wide range of issues instead of focusing on a few specific topics. In summary, of the two hypotheses presented at the beginning of this subsection, the most accurate is that parties tend to have different agendas and rarely promote or debate the same topics.
To further investigate similarities and differences between parties, a machine learning model is built to try to predict the party author of an ad (RQ3.b). For this multi-class classification task, a multinomial logistic regression model is trained on the ads relevant to the 2023 federal elections. Ads are weighted according to their total number of impressions to ensure that ads with a larger reach have a proportionally larger impact on the model’s parameters. For completeness, three different feature combinations are tested: (1) Demographic, representing the percentage of impressions per age and gender bucket; (2) Topic, which are one-hot encoded topic categories; and (3) a combination of Demographic + Topics.
Table 1: AUC ROC score for each feature combination/party
Demographic Topics Demographic
+Topics
SVP 0.748 0.804 0.869
GRÜNE 0.728 0.775 0.857
SP 0.661 0.776 0.816
GLP 0.653 0.736 0.795
FDP 0.685 0.710 0.792
Die Mitte 0.632 0.722 0.759
In general, the model achieves strong discriminative power across all classes and feature combinations, as shown by Table 1 where is reported the One-vs-Rest goodness of fit measured as the area under the ROC curve. Demographic features alone are sufficient to distinguish the parties with good performance. However, Topic features appear to be stronger predictors, and the best-performing model uses the combined set of all features. In the combined model, all parties achieve an AUC ROC score above 0.75, with top-performing classes like SVP and GRÜNE exceeding 0.85.
The recall-normalized confusion matrix (Figure 1(a)) highlights a more complex situation than the model’s high discriminative power might suggest. While recall is high for parties at the poles of the political spectrum, such as the SVP (0.78), it is notably low for the centrist parties Die Mitte and the GLP. This is particularly evident for Die Mitte, which achieves a recall of only 0.08. The confusion matrix shows that Die Mitte ads are most frequently misclassified as SVP (32%) and FDP (24%).
Figure 8: Marginal effects of the model with the best feature combination (demographic + topics). Only coefficients with a p-value below 0.05 are reported.
This result suggests that while the model can reliably distinguish Die Mitte from non-Die Mitte ads (as shown by its high AUC of 0.76), it fails to classify them as the top choice, probably because of a strong feature overlap with other parties. In line with the analyses conducted for RQ1.a and RQ1.b, the difficulty of predicting centrist parties reinforces the observation that they rely less on microtargeting strategies and avoid focusing on specific issues.
As a final step, the marginal effects of the model trained only on Topic features are studied to better understand the predictive interaction between topics and parties. The topic-share plot in Figure 7 represents the raw association between a topic and a party, meaning it also includes the effects of all other topics. In contrast, marginal effects isolate the predictive effect of adding a single topic while holding all others constant, thereby highlighting the influence of each topic. Figure 8 presents the marginal effects that are statistically significant ($p<0.005$), revealing several patterns. For example, environmental and social issues sharply divide the parties. In particular, topics related to “Climate & Environment” are negatively associated with the SVP, the FDP, and Die Mitte, while “Energy Policy” topics are a positive predictor only for Die Mitte. In contrast, “Social Justice & Equality” issues are strong positive predictors of the SP and the GRÜNE. Other policy areas also show distinct alignments. “Healthcare System” topics are positively associated with the SP and Die Mitte, and “Education System” issues appear as relevant predictive topics for the FDP and Die Mitte. Moreover, “Immigration & Asylum” issues are a significant, strong, positive predictor exclusively for the right-wing party SVP.
5 CONCLUSIONS
Switzerland is renowned for its extensive use of direct democracy, which is characterized by frequent referendums and popular votes. Against this backdrop, this paper presents a large-scale, quantitative analysis of more than 40,000 political ads published on Facebook and Instagram in Switzerland. The dataset collected for this study comprises nearly 4.5 million Swiss francs in expenditures and more than 560 million impressions. This is a surprisingly high volume of activity given the relatively small Swiss electorate of 5.6 million.
The findings confirm that major political parties rely heavily on online advertising, using it not only to secure votes in federal elections, but also to influence referendum outcomes. The volume of ads is significantly higher before referenda and greater exposure to pro-Yes advertising has been shown to correlate statistically with referendum approval. In contrast, exposure to pro-No advertising does not show a systematic link to rejection outcomes.
Furthermore, there is significant evidence that political parties widely employ microtargeting based on demographics, such as age and gender. Generally, centrist and right-wing parties tend to attract older male audiences, whereas left-wing parties tend to focus on young women. A more granular analysis reveals that political parties adapt their microtargeting strategies to align also with Switzerland’s distinct linguistic regions. In fact, Switzerland is characterized by deep-seated heterogeneity, not only linguistically but also socio-economically, with significant disparities in wealth, industrial focus, and urban-rural landscapes. As literature has shown, Swiss heterogeneity is a primary driver of political cleavages [OeschRennwald] and voting behavior [SteenbergenMarco, CamatarriFavero], strongly correlating with economic inequalities and urbanity [Mantegazzi03072021]. Microtargeting analysis has several limitations imposed by the Meta Ad Library API, which provides an incomplete picture of targeting mechanisms. First, aside from the opacity of ad selection, the API only reveals targeting parameters related to geographical location, age, and gender. This omits the vast array of other attributes that Meta ad algorithm is known to use for profiling, such as international migration [ZagheniGummadi] or health-related behaviors [MejovaWeberBenevenuto]. Second, the available data does not distinguish between targeting decisions made manually by the campaign advertiser and those driven by Meta’s ad-delivery algorithms.
The content analysis of the ads reveals a clear pattern of issue divergence. Rather than engaging in direct debate with different views on the same issues, parties tend to focus their campaigns on distinct sets of topics, effectively “talking past each other”. This behavior aligns with the well-established political science theory of “issue ownership” [Petrocik], which posits that parties strategically emphasize topics where they have a long-standing reputation for competence [BudgeFarlie1983]. By selectively focusing on “their” issues, parties avoid competing on unfavorable ground and instead attempt to make their own issue priorities the most important for voters [amoros2013issue, BELANGER2008477]. Finally, the results confirm that the audience reach and topics promoted in advertisements differ significantly enough across parties to be used as effective features for a machine learning model to predict an ad’s author.
This study contributes to the ongoing debate about the impact of online political advertising on democratic processes. Although the potential consequences of this phenomenon are still being debated, censoring ads on social media platforms may not be preferable. Instead, transparency of data and algorithms, and data-driven analyzes, would empower users by providing them with the necessary information to understand how they are being targeted.
APPENDIX
Appendix A DATA
Table A1: Thematic Keyword Groups for Data Collection
English Theme German Keywords Italian Keywords French Keywords
To vote / To elect "wählen", "wähle", "wählen wir" "votare", "vota", "votiamo", "voter", "vote", "votons",
"eleggere", "eleggi", "eleggiamo" "élire", "élis", "élisons"
Election(s) / Vote(s) "Wahl", "Wahlen" "elezione", "elezioni", "élection", "élections",
"votazioni", "voto" "votes", "vote"
Poll / Ballot "Abstimmungen", "Abstimmung" (n/a) (n/a)
Referendum "Referendum" "referendum" "référendum"
Electoral Campaign "Wahlkampf", "Wahlkampagne" "campagna elettorale" "campagne électorale"
Electoral Round "Wahlgang" "turno elettorale" "tour électoral"
Counting / Tally "Auszählung" "scrutinio" "scrutin"
Table A2: Impressions by party and language
SP GRÜNE GLP Die Mitte FDP SVP
All 90.0M 55.9M 20.3M 20.3M 49.9M 45.4M
DE 59.4M 37.1M 13.8M 14.7M 27.6M 34.8M
FR 18.0M 15.6M 5.5M 3.8M 19.2M 6.6M
IT 12.6M 3.2M 1.0M 1.8M 3.1M 4.0M
Table A3: Number of ads by party and language
SP GRÜNE GLP Die Mitte FDP SVP
All 3.2K 2.6K 1.4K 1.4K 3.0K 2.5K
DE 2.1K 1.7K 1.1K 1.0K 2.0K 2.1K
FR 0.7K 0.7K 0.2K 0.3K 0.9K 0.2K
IT 0.4K 0.1K 0.1K 0.1K 0.1K 0.2K
Table A4: Expenditure on ads by party and language
SP GRÜNE GLP Die Mitte FDP SVP
All 758.4K 353.1K 157.6K 188.5K 372.7K 353.0K
DE 484.2K 229.1K 113.9K 138.7K 226.1K 286.9K
FR 149.5K 103.5K 37.9K 39.4K 132.2K 39.4K
IT 124.7K 20.5K 5.8K 10.3K 14.4K 26.6K
Appendix B METHODS
Table A5: Results of the Wilcoxon signed-rank test (p < 0.001).
Impressions Expenditure
All 69456 68337
DE 62227 64002
FR 61751 58278
IT 35867 33361
Figure A1: Confusion matrices for the Demographic + Topics model.